OpenStudy (kylav210):

Regan runs and bicycles every day for a total of 60 minutes. Her body uses 9 Calories per minute during running and 7 Calories per minute during bicycling. a) Write and simplify an expression for the total Calories Regan uses running and bicycling each day. b) How many Calories does she use on a day when she runs for 20 minutes?

OpenStudy (anonymous):

1. x=time running; y=time cycling 9x+7y=calories used 2. (20 times 9) + (40 times 7) = 180 + 280 = 460

OpenStudy (anonymous):

i hope it helps;)

OpenStudy (kylav210):

Thanks!!(:

OpenStudy (anonymous):

ur welcome :)